# Quantum Machine Learning for Digital Health? A Systematic Review

### v1

@@ Abstract
With the digitization of health data, the growth of electronic health and medical records lowers barriers for using algorithmic techniques for data analysis. While classical machine learning (ML) techniques for health data approach commercialization and incorporation into clinical workflows, there is not yet clear evidence whether quantum machine learning (QML) will actually provide any empirical advantage for digital health data processing tasks. In this systematic literature review we assess whether, in light of developing digital health technologies, quantum machine learning algorithms have the potential to outperform existing classical methods in efficacy or efficiency. We include digital electronic health/medical records (EH/MRs) and data considered to be a reasonable proxy to EH/MRs. Eligible QML algorithms must be designed for quantum computing hardware, as opposed to quantum-inspired techniques. PubMed, Embase, IEEE, Scopus and a physics preprint server arXiv yielded 4915 studies between 2015 to 10 June 2024. After screening 169 eligible studies, most studies contained widespread technical misconceptions about QML and we excluded 123 studies for insufficient rigor in analysis. Of the remaining 46 studies, only 16 studies consider realistic QML operating conditions, either by testing algorithms on quantum hardware, or using noisy quantum circuits when assessing QML algorithms. We find QML applications in digital health focus primarily on clinical decision support rather than health service delivery or public health. Nearly all QML models are linear quantum models, and therefore represent a subset of general quantum algorithm structures. Meanwhile, novel data-encoding strategies do not address scalibility issues, except in regimes requiring restrictive assumptions about quantum hardware, rendering these protocols inefficient for the general encoding of large health datasets. Our work establishes the current state of evidence for QML-based health applications, paving the way for meaningful QML use-case discovery in digital health.
@@ 1 Introduction
Recent years have seen a proliferation of research proclaiming the utility of quantum machine learning (QML) algorithms for analyzing classical data in many sectors, e.g. finance, cybersecurity, logistics, pharmaceuticals, energy, minerals, and healthcare. With the increasing digitization of health data, the growth of electronic health and medical records [1] paves the way for the use of algorithmic techniques – quantum or classical – for analyzing this data. Potential digital health applications could include clinical decision support, clinical predictive health and health monitoring, public health applications and improving health services delivery and data fusion [2, 3, 4, 5]. The potential for use-case discovery for QML in healthcare [6], biomedical [7] and clinical medicine applications [8] is found to be compelling in previous reviews. However, quantum technologies and classical machine learning are often grouped as interchangeable tools for information processing, creating widespread confusion about regimes in which quantum methods should be applied in lieu of classical methods to solve information processing tasks. Clarifying the algorithmic structure of quantum technologies and how these quantum algorithms are being used in a classical information processing pipeline is thus critical for meaningful dialogue about digital health applications.
@@
Characterizing the role of QML algorithms applied to real-world classical data is nuanced and a challenging question in applications development but also in fundamental QML theory [9, 10]. Quantum advantage refers to asymptotic reduction in computational resources (or some other metric [11]) required by quantum algorithms when compared to classical counterparts, i.e. resources are saved as problem size scales to infinity. Empirical quantum advantage [12] colloquially refers to finite-sized simulations or experiments using quantum over classical algorithms to perform a task, where one assumes any desired resource-savings will scale to larger problems, e.g. in qubit number, high-dimensional or highly structured datasets. However, for classical datasets of arbitrary structure such as those encountered in healthcare settings, there is no known theoretically provable quantum advantage [11]. Instead, the field relies on mostly empirical analysis of QML performance for a variety of pseudo-real-world data, where performance differentials between quantum and classical methods on these smaller problems constitute evidence for testing empirical quantum advantage. In certain mathematical regimes, the role of quantum algorithms for solving inference tasks has been replaced by equivalent classical capability, in a process known as dequantization (e.g. [13, 14]). Meanwhile, most computational analysis of scaling behavior assumes ideal operating conditions and it is unknown if QML methods will retain any benefits in realistic operating settings, such as on near-term noisy quantum hardware.
@@
As typical in medical research settings, a systematic literature review is a standard approach for assessing the strength of evidence for proposed interventions in clinical contexts and public health [15]. In this work, we undertake a systematic literature review of QML applications in digital health between 2015 and 2024. Based on existing evidence in literature, we ask whether quantum machine learning algorithms potentially outperform existing classical methods in efficacy or efficiency for digital health. We analyze the role quantum computing and quantum machine learning algorithms play in processing digital health data. Our objective is to assess the strength of the evidence and dominant trends associated with using QML algorithms for digital health. Our review synthesizes and assesses the current-state evidence including any empirical trends, but by design, the medium of a systematic review cannot speculate on the general future potential for use-case discovery of QML in digital health.
@@
The structure of this document is as follows. We begin by providing an overview of quantum machine learning in Section 2. In Section 3, we outline our approach in accordance with systematic review procedures and guidelines documenting database selection, screening procedures and inclusion criteria for articles. We tabulate key research themes and conduct meta-analysis of our final set of articles in Section 4. Finally, we return to address the research questions of our review, and comment on research opportunities, gaps, limitations and future work in Section 5 before providing concluding remarks in Section 6.
@@ 2 Overview of quantum machine learning (QML)
Quantum computing refers to a broad category of algorithms, for which it is desired that quantum computing hardware will be required to perform some of the computations. Quantum machine learning algorithms are a subset of quantum computation. For the scope of this review, the input of a quantum algorithm is associated with a classical dataset, and an inference problem is defined on the classical dataset. Quantum computational advantage accrues when a quantum algorithm can reduce the number of operations required to solve this inference problem as the size of the problem becomes asymptotically large. Here, the problem size is typically associated with features of the input data e.g. with input data dimension.
@@ 2.2 Algorithms for adiabatic quantum computers
Quantum annealing algorithms assume a very specific type of quantum computing hardware, namely adiabatic computers, (e.g. D-Wave) to solve specific learning tasks. Adiabatic quantum computers can approximately solve computationally hard (i.e. ‘NP-hard’) problems [17] which are very relevant to classical machine learning. The main class of problems encountered in this review relates to quadratic unconstrained binary optimization (QUBO). Examples of QUBO optimization problems include regression, classification, and data compression tasks. Classical, quantum and hybrid annealers can all approximately solve QUBO optimization problems [18], or be used to draw samples from particular types of probability distributions (e.g. Boltzmann distributions) [19].
@@ 3 Methods
Having discussed broad categories of QML algorithms, we now turn to the research question that determines the design of our systematic review. The use of QML algorithms in health settings encompasses a broad range of clinical applications. We use the SPICE framework [30] to ask:
@@
A systematic review was conducted in line with the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) [31] (Appendix B) and was registered on PROSPERO (ID: CRD42024562024) [32]. Screening and data extraction were performed in Covidence [33]. Commonly used nomenclature encountered in this review is summarized in Appendix C.
Figure 3: Search strategy in the SPICE framework [30]. Columns represent digital health setting and perspective (S,P), quantum intervention (I), classical comparator (C), and study characteristics for evidence evaluation (E). Search concepts within columns are combined with logical ‘OR’ statements, while independent columns are concatenated with logical ‘AND’ statements. Search concept performance differs by database e.g. PubMed, Embase (shown above) support reliable MESH explosion of health concepts; while Scopus, arXiv and IEEE instead rely on root word truncation. Complete search syntax modifications by database are enclosed in Appendix D.
Search strategy: Our search strategy is formed by decomposing our research question into elements of the SPICE framework [30], as summarized in Fig. 3. Only articles published after 2015 were included, as the first commercially-available quantum computer was made accessible in 2016 [34] and digitization of health information into electronic records [1] is relatively recent. Hence both factors prohibit meaningful applications development prior to this date. Search syntax was refined by trial and error on PubMed (Fig. 3) in consultation with a health research librarian, and adapted to other databases (Embase, Scopus, arXiv and IEEE, refer Appendix D). Key articles were identified as litmus tests to sense check database-specific search term strategies. Searches were conducted from 10 May to 10 June 2024.
@@
Data extraction and study quality appraisal: Study characteristics were extracted for all included studies. Additionally, a study quality appraisal was performed to form consensus-based decisions about including or excluding particular studies based on robustness [35]. These appraisals are typically implemented during data extraction and prior to narrative synthesis [35]. Our study quality assessment criteria analyses the rigor with which QML algorithms were investigated [9] and we do not include a myriad of other potential benchmarks, e.g. for clinical robustness. At least two reviewers independently scored eligible studies, and the maximum score over both reviewers was selected during consensus formation. Attributes of low vs. high quality studies with were compared with respect to our criteria. Full data extraction template is enclosed in Appendix E and extracted data as well as underlying analysis code for data extraction is available online [36].
@@ 4.1 Characterization of synthesized studies
Our systematic review is summarized by a PRISMA diagram in Fig. 5(a). Our searches identified 4915 distinct studies. A total of 313 studies passed title and abstract screening and went through full-text screening, of which 169 met eligibility criteria. Inter-rater reliability as measured by Cohen’s kappa was substantial for title and abstract screening (0.72) and moderate for full text screening (0.48). According to the distribution of exclusion reasons shown in Fig. 5(b), the most frequent cause for full-text exclusion during screening was distinguishing between genuinely quantum algorithms designed to run on quantum hardware, vs. classical computation invoking ideas, insights or jargon from quantum physics.
@@
Figure 8: Overview of setting, interventions, and comparators for all 16 synthesized studies (columns), including studies with borderline or weak consensus (shaded gray columns). For each study, we indicate primary quantum interventions (top), classical comparators (middle), and digital health datasets (bottom). Quantum interventions deployed on non-digital health datasets are excluded. Interventions or comparators (rows) combine similar multiple-choice data extraction entries. QNNs exclude QCNNs and no VQCs / PQCs were found to contain mid-circuit measurements (MCM) or adaptive gates.
Figure 9: Selected test scores (accuracy, AUC and squared error $\in[0,1]$) reported for quantum interventions deployed on quantum hardware using health data. Ranges reflected min and max test scores attained over different experimental configurations reported in each study; error bars for each test score are shown if originally reported. Synthesized studies with strong (weak) consensus are shaded in green (gray). Ideal scores for accuracy and AUC metrics (unity) differ from squared error metrics (zero). Performance benchmarks that cannot be easily compared or interpreted alongside other metrics, or are not consistently reported, e.g. F1 scores and precision, are excluded for readability. NA: Not applicable. NR: Not explicitly reported or unclear.
The growth in the number of eligible articles on quantum algorithms for digital health seems almost exponential in Fig. 7(a). These applications are broadly categorized into diagnosis, predictive health, and ‘other’ in Fig. 7(c), where an example of the latter could include synthetic health data generation for classical machine learning. Despite this growth, the majority of studies did not enable code and data accessibility.
@@ 4.2 Empirical evidence from synthesized studies
In facilitating a comparison between quantum vs. classical machine learning, Das (2023) [38], Kawaguchi (2023) [39] and Qu (2023) [37] studies are not represented in Fig. 9 due to the lack of reportable metrics for quantum hardware experiments, while Yano (2020) [40] and Landman (2022) [41] are represented in Fig. 9 but do not report a numerical classical benchmark. The remaining 11 studies reported in Fig. 9 reveal major issues in facilitating the comparison between quantum vs. classical interventions. There are three scientifically concerning flaws:
@@ 5 Discussion
Nearly all quantum algorithms were linear quantum models. Some theoretical evidence shows that linear quantum models will require exponentially more qubits than non-linear models [16], and heuristic evidence shows that certain types of linear quantum models will not be useful for the analysis of classical datasets [10]. Even broadening to a larger pool of 169 eligible studies, non-linear quantum models were not encountered. Of our synthesized studies, seven studies used linear quantum kernel methods including Moradi (2022, 2023) [43, 44], Yano (2020) [40], Aswiga (2024) [51], Krunic (2022) [12] and Kawaguchi (2023) [39]. For non-kernel methods, the underlying technologies for Nirula (2021) [52], Qu (2023) [37] and Das (2023) [38] can be recast in linear form. Finally, the quantum transformers and data encoding strategies that yield favourable scaling properties in Cherrat (2024) [48], Landman (2022) [41], and Kazdaghli (2024) [42] use methods developed in Ref. [49]. Aside from a variant proposed in Cherrat (2024), the data encoders and neural networks leveraged by these studies all appear to be described by the framework of linear quantum models. Indeed, the observed absence of clear, consistent performance trends in the empirical meta-analysis of the previous section could in part be explained by the underlying linear quantum models used for many of the studies.
Despite the fact that all quantum models were trained by a supervised learning problem, no study explicitly characterized their optimization landscape. It is well known that optimization of supervised QML algorithms can be plagued by exponentially vanishing gradients (barren plateaus)[53], exponential concentration of kernel values [54], or exponentially concentrated local minima [55]. However only two out of sixteen synthesized studies mentioned optimization challenges associated with their proposed methods for supervised quantum machine learning. Only Cherrat (2024) [48] and Landman (2022) [41] pointed out that their proposed QML methods’ structures may indeed avoid barren plateaus. Whether these circuits avoid exponential concentration remains open. Meanwhile no substantial improvements are found in reducing shot number requirements for QML methods.
@@
We note that this review does not address the future state of QML in health applications. Many open questions remain in the use of quantum algorithms to solve classical inference problems. As examples: whether quantum kernels can offer an advantage in classifying real world data; the role of data vs. quantum model design for attaining quantum advantage, the effects of optimization protocols and noise on overall QML performance, or the existence of quantum sampling or optimization algorithms that could be of use in classical machine learning.

### v2

@@ Abstract
With the digitization of health data, the growth of electronic health and medical records lowers barriers for using algorithmic techniques for data analysis. While classical machine learning (ML) techniques for health data approach commercialization and incorporation into clinical workflows, there is not yet clear evidence whether quantum machine learning (QML) will actually provide any empirical advantage for digital health data processing tasks. In this systematic literature review we assess whether, in light of developing digital health technologies, quantum machine learning algorithms have the potential to outperform existing classical methods in efficacy or efficiency. We include digital electronic health/medical records (EH/MRs) and data considered to be a reasonable proxy to EH/MRs. Eligible QML algorithms must be designed for quantum computing hardware, as opposed to quantum-inspired techniques. PubMed, Embase, IEEE, Scopus and a physics preprint server arXiv yielded 4915 studies between 2015 to 10 June 2024. After screening 169 eligible studies, most studies contained widespread technical misconceptions about QML and we excluded 123 studies for insufficient rigor in analysis. Of the remaining 46 studies, only 16 studies consider realistic QML operating conditions, either by testing algorithms on quantum hardware, or using noisy quantum circuits when assessing QML algorithms. We find QML applications in digital health focus primarily on clinical decision support rather than health service delivery or public health. Nearly all QML models are linear quantum models, and therefore represent a subset of general quantum algorithm structures. Meanwhile, novel data-encoding strategies do not address scalability issues, except in regimes requiring restrictive assumptions about quantum hardware, rendering these protocols inefficient for the general encoding of large health datasets. Our work establishes the current state of evidence for QML-based health applications, paving the way for meaningful QML use-case discovery in digital health.
@@ 1 Introduction
Recent years have seen a proliferation of research proclaiming the utility of quantum machine learning (QML) algorithms for analyzing classical data in many sectors, e.g. finance, cybersecurity, logistics, pharmaceuticals, energy, minerals, and healthcare. With the increasing digitization of health data, the growth of electronic health and medical records [1] paves the way for the use of algorithmic techniques – quantum or classical – for analyzing this data. Potential digital health applications could include clinical decision support, clinical predictive health and health monitoring, public health applications and improving health services delivery and data fusion [2, 3, 4, 5]. The potential for use-case discovery for QML in healthcare [6] and biomedical [7] applications is found to be compelling in previous systematic reviews. Other broader reviews on quantum computing for health, biology and lifesciences [8, 9, 10, 11, 12, 13, 14, 15] hypothesize the potential utility of QML algorithms or quantum subroutines in health, but none of these works are rigorous systematic reviews (and thus reproducible). Indeed, across all of these standard and systematic reviews, we find that the strength of the current evidence base even under mildly realistic operating conditions is not examined.
@@
Characterizing the role of QML algorithms applied to real-world classical data is nuanced and a challenging question in applications development but also in fundamental QML theory [16, 17]. Quantum advantage refers to asymptotic reduction in computational resources (or some other metric [18]) required by quantum algorithms when compared to classical counterparts, i.e. resources are saved as problem size scales to infinity. Empirical quantum advantage [19] colloquially refers to finite-sized simulations or experiments using quantum over classical algorithms to perform a task, where one assumes any desired resource-savings will scale to larger problems, e.g. in qubit number, high-dimensional or highly structured datasets. However, for classical datasets of arbitrary structure such as those encountered in healthcare settings, there is no known theoretically provable quantum advantage [18]. Instead, the field relies on mostly empirical analysis of QML performance for a variety of pseudo-real-world data, where performance differentials between quantum and classical methods on these smaller problems constitute evidence for testing empirical quantum advantage. Most computational analysis of scaling behavior assumes ideal operating conditions and it is unknown if QML methods will retain any benefits in realistic operating settings, such as on near-term noisy quantum hardware. In some cases, the role of quantum algorithms for solving inference tasks has been entirely replaced by equivalent classical capability, in a process known as dequantization (e.g. [20, 21]).
@@
In this work, we undertake a systematic literature review of QML applications in digital health between 2015 and 2024. As typical in medical research settings, a systematic literature review is a standard methodological approach for assessing the strength of evidence for proposed interventions in clinical contexts and public health [22]. Based on existing evidence in literature, we ask whether quantum machine learning algorithms potentially outperform existing classical methods in efficacy or efficiency for digital health. Our objective is to assess the strength of the evidence and dominant trends associated with using QML algorithms for digital health, including assessing the extent to which performance robustness of proposed QML algorithms has been characterized.
@@
The structure of this document is as follows. We begin by providing an overview of quantum machine learning in Section 2. In Section 3, we outline our approach in accordance with systematic review procedures and guidelines documenting database selection, screening procedures and inclusion criteria for articles. We tabulate key research themes and conduct meta-analysis of our final set of articles in Section 4. Finally, we return to address the research questions of our review, and comment on research opportunities, gaps, limitations and future work in Section 5 before providing a future outlook and concluding remarks in Section 6 and Section 7 respectively.
@@ 2 Overview of quantum machine learning (QML)
Quantum computing refers to a broad category of algorithms, for which it is desired that quantum computing hardware will be required to perform some of the computations. Quantum machine learning algorithms are a subset of quantum computation. For the scope of this review, the input of a quantum algorithm is associated with a classical dataset, and an inference problem is defined on the classical dataset.
Quantum computational advantage accrues when a quantum algorithm can reduce the number of operations required to solve this inference problem as the size of the problem becomes asymptotically large. Here, the problem size is typically associated with features of the input data e.g. with input data dimension. From a computer science perspective, algorithms can either improve on the number of queries or samples required (sample complexity) or the number of parallelizable quantum operations (time complexity, or runtime). When quantum algorithms enable improvements in either sample or time complexity, this is sometimes informally referred to as ’quantum advantage’ or ‘speed-up’. An additional metric of memory complexity quantifies the size or type of classical data structures required to store classical inputs, outputs and efficiently recall intermediary classical information during computation, but memory complexity is typically not encountered in the literature for quantum algorithms. A comparison of computational costs of selected classical vs. quantum algorithms for ideal mathematical regimes can be found in Ref. [11], but these were not encountered for real-world health data in our review.
@@ 2.2 Algorithms for adiabatic quantum computers
Quantum annealing algorithms assume a very specific type of quantum computing hardware, namely adiabatic computers, (e.g. D-Wave) to solve specific learning tasks. Adiabatic quantum computers can approximately solve computationally hard (i.e. ‘NP-hard’) problems [24] including approximately solving combinatorial optimization problems. The main class of problems encountered in this review relates to quadratic unconstrained binary optimization (QUBO). Examples of QUBO optimization problems include regression, classification, and data compression tasks. Classical, quantum and hybrid annealers can all approximately solve QUBO optimization problems [25], or be used to draw samples from particular types of probability distributions (e.g. Boltzmann distributions) [26]. While more general forms of adiabatic quantum computing than annealing techniques do exist, we did not encounter any within our included literature, and for this reason have not included a discussion of this form of learning.
@@
The subsections above provide a high-level summary of classes of quantum algorithms that were encountered in this review, but cannot be construed as a comprehensive overview of quantum machine learning (see for example, Refs. [27], [28]). We also note that it is possible to realize quantum annealing tasks on gate-based quantum computers, e.g. Ref. [29], and therefore our classification represents one choice of a non-exclusive method for framing the discussion of QML algorithms. In the next subsection, we consider methods for classical data input to QML algorithms.
@@ 3 Methods
Having discussed broad categories of QML algorithms, we now turn to the research question that determines the design of our systematic review. Unlike most systematic reviews, the use of QML algorithms in health settings encompasses a broad range of clinical applications. We use the SPICE framework [40] to ask:
@@
A systematic review was conducted in line with the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) [41] (Appendix C) and was registered on PROSPERO (ID: CRD42024562024) [42]. Screening and data extraction were performed in Covidence [43]. Commonly used nomenclature encountered in this review is summarized in Appendix D.
Figure 3: Search strategy in the SPICE framework [40]. Columns represent digital health setting and perspective (S,P), quantum intervention (I), classical comparator (C), and study characteristics for evidence evaluation (E). Search concepts within columns are combined with logical ‘OR’ statements, while independent columns are concatenated with logical ‘AND’ statements. Search concept performance differs by database e.g. PubMed, Embase (shown above) support reliable MESH explosion of health concepts; while Scopus, arXiv and IEEE instead rely on root word truncation. Complete search syntax modifications by database are enclosed in Appendix E.
Search strategy: Our search strategy is formed by decomposing our research question into elements of the SPICE framework [40], as summarized in Fig. 3. Only articles published after 2015 were included, as the first commercially-available quantum computer was made accessible in 2016 [44] and digitization of health information into electronic records [1] is relatively recent. Hence both factors prohibit meaningful applications development prior to this date. Search syntax was refined by trial and error on PubMed (Fig. 3) in consultation with a health research librarian, and adapted to other databases (Embase, Scopus, arXiv and IEEE, refer Appendix E). Key articles were identified as litmus tests to sense check database-specific search term strategies. Searches were conducted from 10 May to 10 June 2024.
@@
Data extraction and study quality appraisal: Study characteristics were extracted for all included studies. Additionally, a study quality appraisal was performed to form consensus-based decisions about including or excluding particular studies based on robustness [45]. These appraisals are typically implemented during data extraction and prior to narrative synthesis [45]. Our study quality assessment criteria analyses the rigor with which QML algorithms were investigated [16] and we do not include a myriad of other potential benchmarks, e.g. for clinical robustness. At least two reviewers independently scored eligible studies, and the maximum score over both reviewers was selected during consensus formation. Attributes of low vs. high quality studies with were compared with respect to our criteria. Full data extraction template is enclosed in Appendix F and extracted data as well as underlying analysis code for data extraction is available online [46].
@@ 4.1 Characterization of synthesized studies
Our systematic review is summarized by a PRISMA diagram in Fig. 5(a). Our searches identified 4915 distinct studies. A total of 313 studies passed title and abstract screening and went through full-text screening, of which 169 met eligibility criteria. Inter-rater reliability as measured by Cohen’s kappa was substantial for title and abstract screening (0.72) and moderate for full text screening (0.48). According to the distribution of exclusion reasons shown in Fig. 5(b), the most frequent cause for full-text exclusion during screening was distinguishing between genuinely quantum algorithms designed to run on quantum hardware, vs. classical computation invoking ideas, insights or jargon from quantum physics. If the health setting of physics-centric QML studies is not explicit (e.g. for instance, [47]), then these studies will not be returned in search nor pass title and abstract screening for any systematic review.
@@
Figure 8: Overview of setting, interventions, and comparators for all 16 synthesized studies (columns), including studies with borderline or weak consensus (shaded gray columns). For each study, we indicate primary quantum interventions (top), classical comparators (middle), and digital health datasets (bottom). Quantum interventions deployed on non-digital health datasets are excluded. Interventions or comparators (rows) combine similar multiple-choice data extraction entries. QNNs exclude QCNNs and no VQCs / PQCs were found to contain mid-circuit measurements (MCM) or adaptive gates. A list of links to all sixteen studies, including any available codebases and datasets is compiled in Appendix B.
Figure 9: Selected test scores (accuracy, AUC and squared error $\in[0,1]$) reported for quantum interventions deployed on quantum hardware using health data. Ranges reflected min and max test scores attained over different experimental configurations reported in each study; error bars for each test score are shown if originally reported. Synthesized studies with strong (weak) consensus are shaded in green (gray). Best possible scores for accuracy, fidelity and AUC metrics (unity) differ from squared error metrics (zero). Performance benchmarks that cannot be easily compared or interpreted alongside other metrics, or are not consistently reported, e.g. F1 scores and precision, are excluded for readability. NA: Not applicable. NR: Not reported. Est.: Estimates of numerical values implied from visual/graphical data.
The growth in the number of eligible articles on quantum algorithms for digital health seems almost exponential in Fig. 7(a). These applications are broadly categorized into diagnosis, predictive health, and ‘other’ in Fig. 7(c). ‘Diagnosis’ refers to an application that identifies, characterizes or labels current health data with the aim of supporting a clinical diagnosis e.g. classification of medical images or time varying signals. ‘Predictive’ refers to an application that predicts future health information based on current health data, not necessarily to support formulation of a clinical diagnosis e.g. predicting drug efficacy or disease/risk factors. All remaining applications are group under ‘other‘ e.g. generating synthetic ECG signals based on EHR/EMR data. In Fig. 7(d), we observe that the majority of studies did not enable code and data accessibility which are typically both required to enable tests of reproducibility.
@@ 4.2 Empirical evidence from synthesized studies
We compare quantum vs. classical machine learning by reporting selected metrics from sixteen synthesized studies in the remaining columns of Fig. 9. Of these, five studies do not provide sufficient information to address our review question. Das (2023) [49], Kawaguchi (2023) [50] and Qu (2023) [48] do not report numerical metrics for quantum experiments, while Yano (2020) [51] and Landman (2022) [52] report numerical quantum benchmarks but do not report a numerical classical comparator. The remaining 11 studies reported in Fig. 9 reveal major issues in facilitating the comparison between quantum vs. classical interventions. There are three scientifically concerning flaws:
@@ 5 Discussion
Nearly all quantum algorithms were linear quantum models. Some theoretical evidence shows that linear quantum models will require exponentially more qubits than non-linear models [23], and heuristic evidence shows that certain types of linear quantum models will not be useful for the analysis of classical datasets [17]. Even broadening to a larger pool of 169 eligible studies, non-linear quantum models were not encountered. Of our synthesized studies, seven studies used linear quantum kernel methods including Moradi (2022, 2023) [54, 55], Yano (2020) [51], Aswiga (2024) [62], Krunic (2022) [19] and Kawaguchi (2023) [50]. For non-kernel methods, the underlying technologies for Nirula (2021) [63], Qu (2023) [48] and Das (2023) [49] can be recast in linear form. Finally, the quantum transformers and data encoding strategies that yield favourable scaling properties in Cherrat (2024) [59], Landman (2022) [52], and Kazdaghli (2024) [53] use methods developed in Ref. [60]. Aside from a variant proposed in Cherrat (2024), the data encoders and neural networks leveraged by these studies all appear to be described by the framework of linear quantum models. Indeed, the observed absence of clear, consistent performance trends in the empirical meta-analysis of the previous section could in part be explained by the underlying linear quantum models used for many of the studies. After publishing our pre-print, we were made aware of Ref. [64] as an improvement of quantum methods in Landman 2022 and Cherrat 2024, consisting of a non-linear model in Fig. 1(d). While Ref. [64] fails our inclusion criteria, even its inclusion would not affect the overall conclusions of our review.
Despite the fact that all quantum models were trained by a supervised learning problem, no study explicitly characterized their optimization landscape. It is well known that optimization of supervised QML algorithms can be plagued by exponentially vanishing gradients (barren plateaus)[65], exponential concentration of kernel values [66], or exponentially concentrated local minima [67]. However only two out of sixteen synthesized studies mentioned optimization challenges associated with their proposed methods for supervised quantum machine learning. Here, Cherrat (2024) [59] and Landman (2022) [52] stated that their proposed QML methods’ structures may avoid barren plateaus. All studies failed to provide a systematic characterization of their empirical optimization landscape, and the resources utilized by their chosen optimization protocol in practice. Meanwhile no substantial improvements are found in reducing shot number requirements for QML applications considered in this review.
@@
By design, the medium of a systematic review cannot speculate on the general future potential for use-case discovery of QML in digital health. Many open questions remain in the use of quantum algorithms to solve classical inference problems. As examples: whether quantum kernels can offer an advantage in classifying real world data; the role of data vs. quantum model design for attaining quantum advantage, the effects of optimization protocols and noise on overall QML performance, or the existence of quantum sampling or optimization algorithms that could be of use in classical machine learning. However, we comment on research areas of immediate urgency and interest in the next section.
6 Future outlook
Our review highlights that the language of quantum advantage, empirical quantum utility, speed-up, or resource efficiencies are poorly defined and frequently abused notions in literature. QML applications development could benefit from guidelines on what robust quantum vs. classical comparisons look like. Even leaving aside the issue of how to select the best classical comparator, a comprehensive review on the approaches for benchmarking quantum performance is thus of immediate urgency and interest.
Comparing computational cost improvements enabled by quantum algorithms can be a difficult task. As discussed in Section 2, computational costs are theoretically quantified by sample complexity number of queries) or time complexity (number of sequential operations). As examples, for sample complexity, one must ensure that information contained in each query or sample must be comparable across algorithms. Meanwhile for time complexity, operations contain assumptions about hardware capabilities and these assumptions are not always explicitly stated nor consistent. As elucidated by our discussion on quantum transformers in our review, studies assume that groups of quantum operations can be parallelized. This assumption is not hardware-agnostic: certain quantum operations may be parallelized in some architectures and not in others.
In the absence of theoretical assurances on complexity, we have seen in Fig. 9 how empirical studies use performance metrics such as fidelity or accuracy to argue for the ‘utility’ of quantum algorithms in information processing tasks. We have discussed how arguments of empirical utility or advantage must demonstrate both scalability and robustness of performance. Since simulating 100 qubits can be within reach of classical computers, we argue that characterizing properties of QML algorithms as a function of system size is more important than reporting any single figure of merit at some arbitrary choice of system size. Secondly, relying solely on ideal simulations offer no insight into robustness, and one simple test of robustness is to understand and mitigate the impact of noise. Thirdly, we find that the choice of performance metrics in empirical studies is diverse, often ad-hoc, and limits how to perform meta-analysis of evidence in the field.
Figure 10: Proposed minimal requirements for future QML study design, complementing the quality assessment framework used in our review (c.f. Fig. 6).
We summarize these and other considerations to specify minimum requirements for the robust analysis of quantum algorithms on classical datasets. Our proposed qualitative framework is presented in Fig. 10. This framework is complementary to the quality analysis framework used in this review. In our framework, the minimum requirements outlined in Fig. 10 (column 3) can be immediately met. However, many of the ideal requirements in Fig. 10 (column 4) may require new research due to the nascency of the field. Indeed, the field has made some progress towards these ideal requirements: challenges of technical reproducibility in QML are discussed in Ref. [16] and quantum algorithm performance has been linked closely with hardware benchmarking in Ref. [69]. However, research gaps continue to exist, for example, in the lack of a principled approach to link quantum algorithms to structures in classical data; to select appropriate performance metrics; to use noise characterization tools to set performance expectations and benchmark algorithm performance on hardware, and/or to compare quantum methods with a repository of best-in-class classical benchmarks for industry subdomains. Further development and testing of the proposed framework is both urgent and important to urge better empirical evidence in our nascent field. Finally, we note that our review methodology is future-proof: by changing only the search period, we may provide a systematic update on the quality of research evidence for QML in digital health in future.